OpenStudy (anonymous):

Let V = R^2 (all ordered pairs of real numbers) with addition defined in the usual way by (x1+y1) + (x2,y2) = (x1+x2, y1+y2) but with scalar multiplication defined by alpha (x,y) = (alpha x, y). Show that V is not a vector space by demonstrating how one of its axioms fail to hold.

OpenStudy (anonymous):

alpha x hmmmm been a while is alpha, a constant?

OpenStudy (anonymous):

alpha is a scalar

OpenStudy (anonymous):

alright so you're going to have to go through each axiom, i know it's going to be one of the multiplication axioms i believe, let me find the list so i can go through them aso

OpenStudy (anonymous):

on my exam, I thought A5 didn't hold which was alpha (x+y) = alpha x + alpha y for each scalar alpha alpha and any x and y in V but it turns out... its A6 that doesn't hold: (alpha + beta) x = alpha x + beta x for any scalars alpha and beta and any x in V. I just don't understand the reason.

OpenStudy (anonymous):

so for multiplication you have \[u=(\alpha x, y)\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[cu+ku=c(x,y)+k(x,y)=(cx,y)+(kx,y)=(cx+kx,2y)\]

OpenStudy (anonymous):

is that A5 or A6?

OpenStudy (anonymous):

so A5 is c(x,y) = c(x1,x2) + c(y1,y2) = (cx1,x2)+(cy1,y2) = c((x1,y1)+(x2,y2))